INAF LBC+MODS+LUCI queue observing, 2012 October 20-21

Observer: M. Faccini, E. Sani, K. Boutsia
Telescope Operator: D. Gonzalez Huerta
Support Astronomer: M. Edwars
Instrument: MODS

Summary

Clear sky at sunset and winds between 5 - 10 m/s
We start with Lardo M2 project
At 06:24UT we have to CLOSE DOME for High humidity (90%) and we stay closed for the night.
At the end of night we obtained luci darks for yesterday's data.

ATTENTION : MODS calibration obtained tonight were with the telescope at the HORIZON - we will repeat them tomorrow Taking calibrations at the horizon, gave a lot of problems with the laser lock-on on the red side.

Data Summary

Project_RUN Objects Filter ExpTime Notes
Lardo_5B M2 RGB blue_grating 1hr target completed - seeing within specs
Lardo_5B M2 SGB blue_grating 45min repeated 1exp because the first night there were clouds during the second exposure - COMPLETED
Standard bd284211 blue_grating    
Gilli C 2XMMJ22 red_grating 45min we had to close dome for humidity after the 3rd exposure
Standard feige red_grating    

Details

01:27 acqMODS bd284211.acq==mods1r.20121020.0002-mods1r.20121021.0001==
  • seeing 1.2"/1.3" on the guider
  • modsAlign -r mods1r.20121020.0002.fits
  • offset manually -0.1 in X

01:44 execMODS -b BlueGrating bd284211.obs mods1b.20121021.0001-0003
  • seeing 1.2"-1.3"

02:09 Starting with Lardo M2 RGB (ID541705)
  • acqMODS mods.3.m2rgbNew.acq mods1b 0004-0006
  • modsAlign mods.3.m2rgb.mms mods1b.20121021.0004.fits mods1b.20121021.0005.fits

02:16 execMODS mods.3.m2rgb.obs mods1b 0007-0009
  • at the end of 1st frame exposure seeing ~1.0"
  • at the end of 2nd exposure seeing is getting a bit better: 0.9"

03:24 acqMODS mods.4.m2SGBNew.acq mods1b 0010-0012
  • seeing 0.9. repate 1exposure of 3 cause cloudy weather preavious night (prob. out og specs)
  • modsAlign mods.4.m2SGB.mms mods1b.20121021.0010.fits mods1b.20121021.0011.fits

03:39 execMODS mods.4.m2SGBshort.obs mods1b 0013
  • variable seeing between 0.9 and 1.2
  • 4:29 end of project Lardo M2

04:32 Slewing to Gilli C 2XMMJ22 22:35:32.08 +33:57:55.1
  • acqMODS 2xmmj22.acq mods1r 0003-0004
  • Target near the guiding prob, it cover a quarter of the field
  • modsAlign -l mods1r.20121021.0003.fits mods1r.20121021.0004.fits

05:00 guiding probe followed the modalign, we change guider star
  • repeat acqMODS 2xmmj22.acq mods1r 0006-0008
  • modsAlign -l mods1r.20121021.0006.fits mods1r.20121021.0007.fits

05:15 execMODS 2xmmj22.obs mods1r.20121021.0009-0011
  • seeing 0.8
  • humidity reaches 86%, risk to close.........

06:06 We aborted during the last exposure due to high humidity.
  • we slew for the standard
  • acqMODS standard.for.2xmmj22.feige110.acq mods1r 0012-0013
  • modalign -r mods1r.20121021.0012.fits

06:14 execMODS standard.for.2xmmj22.feige110.obs mods1r 0014-0016

06:24 closing dome due to high humidity

07:46 we start MODS calib - since we are closed
  • execMODS mods.3.m2rgb.cal mods1b 14-17
  • execMODS mods.3.m2rgbflatsONLY.cal mods1b 18-28 some flats are SATURATED
  • TEST execMODS mods.3.m2rgbflatONLY.cal (only 1 frame) mods1b 29-32 QTH+ND1 is fine BUT QTH+UG5 some slits are saturated / with exptime 2sec
  • with exptime 2sec they seem to be fine - execMODS mods.3.m2rgbUG5.cal mods1b 33-37
  • verify that the mods.4 flats were not saturated - execMODS mods.4.m2SGBflatONLY.cal mods1b 38-39 They are both OK - we do not need to repeat.
  • execMODS grlamps_bLS1.0.cal mods1b 40-42 arcs blue grating with 1.0"slit
  • execMODS grlamps_r.cal mods1r 0017-0018 abort because Xe+Kr on while we were supposed to be doing Ne NO IMCS lock
  • execMODS grpixflats_r.cal mods1r 0019-0023 NO IMCS lock - tests to fix this - we turn MODS off and ON
  • execMODS grlamps_r.cal mods1r 0025-0028 repeat - it skips the Ne exp
  • execMODS grlamps_rNe.cal mods1r 0030 it did not work ... the lamp is OFF
  • execMODS grlamps_rNe.cal mods1r 0031-0032 Ne arcs
  • execMODS SlitFlatLBT5_r.cal mods1r 0033-0037
  • execMODS grflats_r.cal mods1r 0038-042 SATURATED!
  • execMODS grflats_r1.0.cal mods1r 0043-047 the first saturated and the rest not
  • execMODS grpixflats_b.cal mods1b 0044-0?? repeat the ones we took this afternoon - the first block had the yesterday's UT date mods1b.20121020.0015-0019

LUCI darks

-- RobertoSpeziali - 21 Oct 2012
Topic revision: r7 - 14 Jan 2013, MarcoFaccini
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback